OpenStudy (anonymous):

2. Tap A fill tank in 6mins, Tap B fill in 9mins. Pipe C empty the tank in 15mins, what is the time taken to fill up the tank when the 2 taps and pipe are in the tank ?

OpenStudy (anonymous):

da/dt = 6 db/dt = 9 dc/dt = -15

OpenStudy (anonymous):

am i heading in the right direction ?

OpenStudy (mathmath333):

is the answer \(4.73~min\) i need to confirm

OpenStudy (mathmath333):

the part filled in 1 min =\(\dfrac{1}{9}+\dfrac{1}{6}-\dfrac{1}{15}=\dfrac{19}{90}\) and the time taken to fill it is \(\dfrac{90}{19}\)

OpenStudy (mathmath333):

so its equal to 4 min 44 sec

OpenStudy (anonymous):

please @HatcrewS if you kno chinese could you please help meh there is no one in the chinese subject.